OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –1 6 0 4 Line 2 x y 0 7 1 5 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

Use point-slope form to write an equation for each line. line 1 m = (4-6)/(0--1) = -2 y - (4) = -2 (x - (0)) y - 4 = -2x y = -2x + 4 line 2 m = (5-7)/(1-0) = -2 y - (7) = -2(x - (0)) y - 7 = -2x y = -2x + 7 Since both lines have the same slope (-2) they are parallel and thus do not intersect. So, there are no solutions.